# Deep learning methods for inverse problems using connections between proximal operators and Hamilton–Jacobi equations

### v1

@@ 4 Learning priors and the inverse problem for Hamilton–Jacobi Equations with incomplete information
In this section, we consider the inverse problem of learning the prior in the proximal operator Eq. 2 with incomplete information: given $t>0$ and a set of samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$, estimate the prior $J$ that best recovers $\bm{x}\mapsto S(\bm{x},t)$. Recall from Theorem 3.1 that when $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous, $\bm{x}\mapsto S(\bm{x},t)$ is reachable if and only if it is semiconcave. In this case, the prior $\bm{x}\mapsto J_{\text{BVS}}(\bm{x})$ obtained from the HJ terminal value problem Eq. 14 provides a prior function that recovers $(\bm{x},t)\mapsto S(\bm{x},t)$ exactly. Hence we will focus on studying how to approximate the prior $J_{\text{BVS}}$ from a set of samples.
Note that if the triplet $(\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t))$ is known, then (i) the function is $\bm{x}\mapsto S(\bm{x},t)$ is differentiable at $\bm{x}$ and (ii) the unique minimum in the Lax–Oleinik formula Eq. 10 can be represented via Eq. 11:
(17) $S(\bm{x}_{k},t)=\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}+J(\bm{y}_{k}),\,\text{with}\ \bm{y}_{k}=\bm{x}_{k}-t\nabla_{\bm{x}}S(\bm{x}_{k},t).$
Moreover, Theorem 3.5 and formula Eq. 6 imply $J(\bm{y}_{k})=J_{\text{BVS}}(\bm{y}_{k})$, $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is convex. Thus one possible approach for estimating $J_{\text{BVS}}$ is to approximate $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ piecewise from below at the points $\left\{\bm{y}_{k}\right\}_{k=1}^{K}$.
We consider the problem of approximating $J_{\text{BVS}}$ piecewise from below and its implications in Section 4.1. This approximation problem turns out to be related closely to max-plus algebra theory for approximating solutions to HJ PDEs [akian2006max, fleming2000max, gaubert2011curse]; we discuss this in Section 4.2. We then consider in Section 4.3 the more general problem of learning a convex function to approximate $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ directly, applying the discussions in Section 4.1Section 4.2.
@@ 4.1.1 Piecewise affine approximation
We first consider the PAM approximation of the convex function $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$
@@
$tJ_{\text{PAM}}(\bm{y})=\max_{k\in\{1,\dots,K\}}\left\{tJ_{\text{BVS}}(\bm{y}_{k})+\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}\right\}.$
How good is $J_{\text{PAM}}$ as initial condition for the HJ PDE Eq. 3? In light of Theorem 3.5, $J_{\text{PAM}}$, unsurprisingly, cannot reconstruct $\bm{x}\mapsto S(\bm{x},t)$. Indeed, a formal calculation yields
(19) $\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{PAM}}(\bm{y})\right\}=\begin{cases}S(\bm{x}_{k},t)&\ \text{if $\bm{x}=\bm{x}_{k}$, $k\in\{1,\dots,K\}$},\\ +\infty,&\ \text{otherwise}.\end{cases}$
See Section A.1 for details. Thus approximating $J_{\text{BVS}}$ via its PAM approximation recovers the samples $\{S(\bm{x}_{k},t)\}_{k=1}^{K}$ but nothing else.
@@ 4.1.2 Piecewise quadratic approximation
Here, we assume $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})$ is semiconvex with constant $1-\alpha$ with $\alpha>0$, so that $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is $1-\alpha$ strongly convex. We can then approximate this strongly convex function via its PQMs:
@@
(20) $tJ_{\text{PQM}}(\bm{y})=\max_{k\in\{1,\dots,K\}}\left\{J(\bm{y}_{k})+\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}+\frac{\alpha}{2}\left\|{\bm{y}-\bm{y}_{k}}\right\|_{2}^{2}\right\}.$
How good is $J_{\text{PQM}}$ as an initial condition for the HJ PDE Eq. 3? Again, in light of Theorem 3.5, $J_{\text{PQM}}$ cannot reconstruct $\bm{x}\mapsto S(\bm{x},t)$. Nonetheless, a formal calculation yields
@@
for some $k\in\{1,\dots,K\}$. See Section A.2 for more details. Hence $J_{\text{PQM}}$ leads to an approximation of $(\bm{x},t)\mapsto S(\bm{x},t)$ that is finite everywhere. In the next section, we describe how max-plus algebra theory [akian2006max, fleming2000max, gaubert2011curse] can be used to quantify the approximation errors more precisely.
@@ 4.2 Max-plus algebra theory for Hamilton–Jacobi PDEs and approximation results
We consider here max-plus algebra techniques for approximating solutions to certain HJ PDEs. Let $\alpha>0$ and let $\Psi\colon\mathbb{R}^{n}\to\mathbb{R}$ denote a $(1-\alpha)$-semiconvex function obtained. Following [gaubert2011curse, Section III], we approximate $\Psi$ using $K$ vectors $\{\bm{p}_{k}\}_{k=1}^{K}\subset\mathbb{R}^{n}$ with $K$ semiconvex functions $\bm{y}\mapsto\langle\bm{p}_{k},\bm{y}\rangle-\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ and a function $a\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$:
@@
Here, we suppose the vectors $\{\bm{p}_{k}\}_{k=1}^{K}$ and $\bm{p}\mapsto a(\bm{p})$ are selected so that $\Psi_{\text{MP}}(\bm{y})\leqslant\Psi(\bm{y})$. As discussed in Section 4.1, such a selection is possible via the affine piecewise quadratic minorants of the $(1-\alpha)$-strongly convex function $\bm{y}\mapsto\Psi(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$. Let $\mathcal{Y}$ denote a full dimensional compact, convex subset of $\mathbb{R}^{n}$ and consider the $L_{\infty}$ error
@@
The following result from max-plus algebra theory, proven in [gaubert2011curse], stipulates that whatever vectors $\{\bm{p}_{k}\}_{k=1}^{K}$ and function $\bm{p}\mapsto a(\bm{p})$ are used to approximate $\Psi$, the minimal $L_{\infty}$ error scales as an inverse power law in $K$ and the dimension $n$ in the limit $K\to+\infty$.
@@
Let $\alpha>0$, and let $\mathcal{Y}$ denote a full-dimensional compact, convex subset of $\mathbb{R}^{n}$. If $\Psi\colon\mathbb{R}^{n}\to\mathbb{R}$ is twice continuously differentiable and $1-\alpha$ semiconvex, then there exists a constant $\beta(n)>0$ depending only on $n$ such that
(23) $\delta_{\infty}(\Psi,K,\mathcal{Y})\sim\beta(n)\left(\frac{1}{K}\int_{\mathcal{Y}}(\det\left(\nabla_{\bm{y}}^{2}\Psi(\bm{y})+\bm{I}_{n\times n})\right)^{\frac{1}{2}}\mathop{}\!d\bm{y}\right)^{2/n}$
as $K\to+\infty$.
@@ 4.3 Applications to the inverse problem for Hamilton–Jacobi Equations
We consider here the problem of quantifying approximations of the prior function $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})$ when the latter is sufficiently regularized and when we have access to the values $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$. Max-plus algebra theory provides us with a first approximation result:
Corollary 4.2.
Let $t>0$ and assume $tJ_{\text{BVS}}$ is twice continuously differentiable and $(1-\alpha)$-semiconvex with $\alpha>0$. Let $\mathcal{Y}$ denote a full-dimensional compact, convex set of $\mathbb{R}^{n}$. Then there exists a constant $\beta(n)$ depending only on $n$ such that
(24) $\delta_{\infty}(tJ_{\text{BVS}},K,\mathcal{Y})\sim\beta(n)\left(\frac{1}{K}\int_{\mathcal{Y}}\det\left(t\nabla_{\bm{y}}^{2}J_{\text{BVS}}(\bm{y})+\bm{I}_{n\times n}\right)^{\frac{1}{2}}\mathop{}\!d\bm{y}\right)^{2/n}$
as $K\to+\infty$.
@@
Immediate from Theorem 4.1 because $J_{\text{BVS}}$ satisfies all its assumptions.
Corollary 4.2 provides a lower bound for the approximation error of $J_{\text{BVS}}$ relative to $J_{\text{PQM}}$. Indeed, Theorem 4.1 and Corollary 4.2 and the fact that $J_{\text{PQM}}(\bm{y})\leqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$ imply
(25) $\delta_{\infty}(tJ_{\text{BVS}},K,\mathcal{Y})\leqslant t\sup_{\bm{y}\in\mathcal{Y}}|J_{\text{BVS}}(\bm{y})-J_{\text{PQM}}(\bm{y})|.$
Thus in this case $J_{\text{PQM}}$ approximates $J_{\text{BVS}}$ from below in $\Omega(1/K^{n/2})$ as $K\to+\infty$. We show below a similar upper bound holds using any reachable function $\tilde{J}\in I_{t}(S)$.
Theorem 4.4.
Let $t>0$ and assume $tJ_{\text{BVS}}$ is twice continuously differentiable and $(1-\alpha)$-semiconvex with $\alpha>0$. Let $\mathcal{Y}$ denote a full-dimensional compact, convex set of $\mathbb{R}^{n}$ and let $\tilde{J}\in I_{t}(S)$ denote a function that can can reach $\bm{x}\mapsto S(\bm{x},t)$. Then
Proof 4.5.
First, note Theorem 3.5 implies $\tilde{J}(\bm{y})\geqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, with equality for every $\bm{y}\in\mathbb{R}^{n}$ for which $\bm{y}=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$ for some $\bm{x}\in\mathbb{R}^{n}$. Thus
@@
Finally, since $J_{\text{BVS}}$ is twice continuously differentiable and $(1-\alpha)$ semiconvex with $\alpha>0$, we can invoke Theorem 4.1 with $\Psi\equiv J_{\text{BVS}}$ to get
$\delta_{\infty}(J_{\text{BVS}},K,\mathcal{Y})\leqslant t\sup_{\bm{y}\in\mathcal{Y}}|\tilde{J}(\bm{y})-J_{\text{PQM}}(\bm{y})|,$
that is, inequality Eq. 26 holds. This concludes the proof.
Theorem 4.4 suggests it is possible to learn $J_{\text{BVS}}$ via a function $\tilde{J}$ that is twice continuously differentiable and (1-$\alpha$)-semiconvex and assess the approximation error using the right-hand-side Eq. 26 as a proxy, in particular by driving $\sup_{\bm{y}\in\mathcal{Y}}|\tilde{J}(\bm{y})-J_{\text{PQM}}(\bm{y})|$ to zero using sufficiently large enough data by training $\tilde{J}(\bm{y})$ appropriately.
In the next section, we consider the problem of learning this function using deep neural networks, specifically learned proximal networks [fang2024whats], to enforce the semiconvexity property required for $\tilde{J}$.

### v2

@@ 4 Learning priors and the inverse problem for Hamilton–Jacobi Equations with incomplete information
In this section, we consider the inverse problem associated to the HJ initial value problem Eq. 3 with incomplete data: given $t>0$ and samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$, learn the prior $J$ that “best” recovers the observations $\{S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$. Assuming $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous, recall from Theorem 3.1 that $\bm{x}\mapsto S(\bm{x},t)$ is reachable if and only if it is semiconcave. If reachable, the backward viscosity solution $\bm{x}\mapsto J_{\text{BVS}}(\bm{x})$ of the HJ terminal value problem Eq. 14 then recovers $(\bm{x},t)\mapsto S(\bm{x},t)$. Thus, we will focus on studying how to learn $J_{\text{BVS}}$.
As a first step, note that the unique minimum in the Lax–Oleinik formula Eq. 10 at the sample $\bm{x}=\bm{x}_{k}$ can be represented via Eq. 11:
(19) $S(\bm{x}_{k},t)=\frac{1}{2t}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}+J(\bm{y}_{k}),\ \text{where}\ \bm{y}_{k}=\bm{x}_{k}-t\nabla_{\bm{x}}S(\bm{x}_{k},t).$
Moreover, Theorem 3.5 implies $J(\bm{y})\geqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, and in particular we have $J(\bm{y}_{k})=J_{\text{BVS}}(\bm{y}_{k})$. Taken together, this gives samples $\{\bm{y}_{k},J(\bm{y}_{k})\}_{k=1}^{K}$ of the true underlying prior $J$ (and $J_{\text{BVS}}$).
As a second step, recall the function $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is convex by Theorem Theorem 3.3. Thus instead of learning $J_{\text{BVS}}$, we can equivalently learn the convex function $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ using the samples $\{\bm{y}_{k},J(\bm{y}_{k})\}_{k=1}^{K}$. Before we propose methods for learning this convex function numerically using deep learning and convex neural networks (the topic of Section 5), we discuss below different approaches for learning $J_{\text{BVS}}$, in particular by approximating $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ from below as a piecewise convex function at the points $\left\{\bm{y}_{k}\right\}_{k=1}^{K}$. We consider this below in Section 4.1. This approximation problem, it turns out, is related to max-plus algebra theory for approximating solutions to HJ PDEs [akian2006max, fleming2000max, gaubert2011curse]; we discuss this in Section 4.2. We specialize these results to the problem of approximating $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ in general in Section 4.3.
@@ 4.1.1 Piecewise affine approximation
We first consider the PAM approximation of the convex function $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$:
@@
(21) $tJ_{\text{PAM}}(\bm{y})=\max_{k\in\{1,\dots,K\}}\left\{tJ_{\text{BVS}}(\bm{y}_{k})+\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}\right\}.$
How good is $J_{\text{PAM}}$ as an initial condition for the HJ PDE Eq. 3? A short calculation yields
(22) $\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+J_{\text{PAM}}(\bm{y})\right\}=\begin{cases}S(\bm{x}_{k},t)&\ \text{if $\bm{x}=\bm{x}_{k}$, $k\in\{1,\dots,K\}$},\\ +\infty,&\ \text{otherwise},\end{cases}$
The PAM approximation recovers $\{S(\bm{x}_{k},t)\}_{k=1}^{K}$ and nothing else. It is therefore generally inadequate to use as a meaningful learned prior or estimate of $J_{\text{BVS}}$.
@@ 4.1.2 Piecewise quadratic approximation
Assume now $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})$ is semiconvex with constant $(1-\alpha)$, $\alpha>0$. Then $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is $(1-\alpha)$ strongly convex, and we can approximate it via its piecewise quadratic minorant (PQM):
@@
$tJ_{\text{PQM}}(\bm{y})=\max_{k\in\{1,\dots,K\}}\left\{J(\bm{y}_{k})+\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}_{k}}\right\|_{2}^{2}-\frac{1}{2}\left\|{\bm{x}_{k}-\bm{y}}\right\|_{2}^{2}+\frac{\alpha}{2}\left\|{\bm{y}-\bm{y}_{k}}\right\|_{2}^{2}\right\}.$
How good is $J_{\text{PQM}}$ as an initial condition for the HJ PDE Eq. 3? A formal calculation, relegated to Appendix A, yields
@@
for some $k\in\{1,\dots,K\}$. The PQM approximation now leads to an approximation of $(\bm{x},t)\mapsto S(\bm{x},t)$ that recovers $\{S(\bm{x}_{k},t)\}$ and is finite everywhere. How accurate it is, however? In the next section, we describe how max-plus algebra theory [akian2006max, fleming2000max, gaubert2011curse] can be used to quantify the approximation errors more precisely.
@@ 4.2 Max-plus algebra theory for Hamilton–Jacobi PDEs and approximation results
We consider here max-plus algebra techniques for approximating solutions to certain HJ PDEs. Let $\alpha>0$ and let $\Psi\colon\mathbb{R}^{n}\to\mathbb{R}$ denote a $(1-\alpha)$-semiconvex function. Following [gaubert2011curse, Section III], we approximate $\Psi$ using $K$ vectors $\{\bm{p}_{k}\}_{k=1}^{K}\subset\mathbb{R}^{n}$ with $K$ semiconvex functions $\bm{y}\mapsto\langle\bm{p}_{k},\bm{y}\rangle-\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ and a function $a\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$:
@@
Here, we suppose the vectors $\{\bm{p}_{k}\}_{k=1}^{K}$ and $\bm{p}\mapsto a(\bm{p})$ are selected so that $\Psi(\bm{y})\geqslant\Psi_{\text{MP}}(\bm{y})$. As discussed in Section 4.1, such a selection is possible via the PQM of the $(1-\alpha)$-strongly convex function $\bm{y}\mapsto\Psi(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$. Let $\mathcal{Y}$ denote a full dimensional compact, convex subset of $\mathbb{R}^{n}$ and consider the $L_{\infty}$ error
@@
The following result from [gaubert2011curse], stipulates that whatever vectors $\{\bm{p}_{k}\}_{k=1}^{K}$ and function $\bm{p}\mapsto a(\bm{p})$ are used to approximate $\Psi$, the minimal $L_{\infty}$ error scales as an inverse power law in $K$ and the dimension $n$ in the limit $K\to+\infty$.
@@
Let $\alpha>0$, and let $\mathcal{Y}$ denote a full-dimensional compact, convex subset of $\mathbb{R}^{n}$. If $\Psi\colon\mathbb{R}^{n}\to\mathbb{R}$ is twice continuously differentiable and $(1-\alpha)$ semiconvex, then there is a constant $\beta(n)>0$ such that
(25) $\delta_{\infty}(\Psi,K,\mathcal{Y})\sim\beta(n)\left(\frac{1}{K}\int_{\mathcal{Y}}\left(\det\left(\nabla_{\bm{y}}^{2}\Psi(\bm{y})+\bm{I}_{n\times n}\right)\right)^{\frac{1}{2}}\mathop{}\!d\bm{y}\right)^{2/n}\ \text{as $K\to+\infty$.}$
@@ 4.3 Applications to the inverse problem for Hamilton–Jacobi Equations
We now return to the setting of Section 4.1, namely to approximate the backward viscosity solution $\bm{x}\mapsto J_{\text{BVS}}(\bm{x})$ using samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$ and assuming, additionally, that $J_{\text{BVS}}$ is sufficiently “regularized”. Max-plus algebra theory provides us with the following approximation result in the “large sample size” setting.
Theorem 4.2.
Let $t>0$ and assume $tJ_{\text{BVS}}$ is twice continuously differentiable and $(1-\alpha)$-semiconvex with $\alpha>0$. Let $\mathcal{Y}$ denote a full-dimensional compact, convex set of $\mathbb{R}^{n}$ and let $\tilde{J}\in I_{t}(S)$ denote a function that can reach $\bm{x}\mapsto S(\bm{x},t)$. Then there exists a constant $\beta(n)$ depending only on $n$ such that
(26) $\delta_{\infty}(tJ_{\text{BVS}},K,\mathcal{Y})\sim\beta(n)\left(\frac{1}{K}\int_{\mathcal{Y}}\left(\det\left(t\nabla_{\bm{y}}^{2}J_{\text{BVS}}(\bm{y})+\bm{I}_{n\times n}\right)\right)^{\frac{1}{2}}\mathop{}\!d\bm{y}\right)^{2/n}$
as $K\to+\infty$, and
@@
The asymptotic limit Eq. 26 follows immediately from Theorem 4.1 because $J_{\text{BVS}}$ satisfies all its assumptions. For inequality Eq. 27, first note Theorem 3.5 implies $\tilde{J}(\bm{y})\geqslant J_{\text{BVS}}(\bm{y})$ for every $\bm{y}\in\mathbb{R}^{n}$, with equality for every $\bm{y}\in\mathbb{R}^{n}$ for which $\bm{y}=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$ for some $\bm{x}\in\mathbb{R}^{n}$. Thus
@@
Finally, since $J_{\text{BVS}}$ is twice continuously differentiable and $(1-\alpha)$ semiconvex with $\alpha>0$, we can invoke Theorem 4.1 with $\Psi\equiv J_{\text{BVS}}$ to get inequality Eq. 27. This concludes the proof.
Thus $J_{\text{PQM}}$ approximates $J_{\text{BVS}}$ from below in $\Omega(1/K^{n/2})$ as $K\to+\infty$, and the $L_{\infty}$ error of any reachable function $\tilde{J}\in I_{t}(S)$ with respect to $J_{\text{PQM}}$ is bounded from below by this minimal error $\delta_{\infty}(tJ_{\text{BVS}},K,\mathcal{Y})$. Thus if there are “enough” samples approximating “well” $\bm{x}\mapsto S(\bm{x},t)$, then the PQM approximation will converge. In addition, its error rate may be small if the determinant term in Eq. 26 is close to zero.
What can be said when $tJ_{\text{BVS}}$ is not $(1-\alpha)$-semiconvex? Without further information on the prior $J$ or other regularizing properties of $tJ_{\text{BVS}}$, the answer is unclear, at least to the best of the knowledge of the authors. Deep learning methods can provide a way to train a meaningful prior to approximate $tJ_{\text{BVS}}$, e.g., by fitting well or perfectly a convex neural network $\bm{y}\mapsto G_{\theta}(\bm{y})$ to the function $\bm{y}\mapsto tJ_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ and enforcing enough regularity [^2] It is enough, e.g., to enforce $\bm{y}\mapsto G_{\theta}(\bm{y})/t-\frac{1}{2t}\left\|{\bm{y}}\right\|_{2}^{2}$ to be uniformly Lipschitz continuous, since then the infimum problem would correspond to the Lax–Oleinik formula to the solution of an appropriate forward HJ PDE [bardi1998hopf, crandall1992user, crandall1983viscosity]. to ensure
$\inf_{\bm{y}\in\mathbb{R}^{n}}\left\{\frac{1}{2t}\left\|{\bm{x}-\bm{y}}\right\|_{2}^{2}+\frac{1}{t}\left(G_{\theta}(\bm{y})-\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}\right)\right\}$
is finite for every $\bm{x}\in\mathbb{R}^{n}$. In the next section, we present some numerical results to test this proposed deep neural network approach.